40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 180 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  7918   Fri Jan 18 12:08:08 2013 KojiUpdateLockingDust?

No

Quote:

My question is, could a few scattering centers cause the behavior that we are seeing?

 

  7917   Fri Jan 18 09:54:18 2013 JenneUpdateLockingPRM - Flat mirror cavity

Quote:

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

 Dang it.  I didn't confirm that the movie was good, just that it was there.  It's corrupted or something, and won't play.  I'll just have to make a new movie today after I realign the cavity.

  7916   Fri Jan 18 00:41:34 2013 JenneUpdateLockingDust?

I was thinking tonight about more possible reasons that our PRC sucks, and I wonder if dust on the BS could create the problem.

Historically, Kiwamu and I found a few dust particle scattering centers every time we inspected the test masses before drag wiping. Sometimes, there would be one frustratingly close to the center of the optic. I'm not sure if we ever made note of how many we saw and where they were, except out loud to the assembled crowd.

Anyhow, the BS is the only IFO optic that was not replaced, so I'm not sure how long it has been since it was cleaned. If the PR-flat cavity looks okay and we take out the BS to do a PRM-ITMY cavity, we should inspect the beam splitter.

Also, the PRM could need cleaning, but at least it has been drag wiped within recent memory.

My question is, could a few scattering centers cause the behavior that we are seeing?

 

EDIT:  List o' elogs....

 

Elog 5301 - Some details on dust seen on ITMs and ETMs, Aug 2011.

Elog 4084 - Kiwamu's in-situ drag wiping how-to, with details on some of the dust we saw. Dec 2010.

Elog 3736 - PRM drag wiped before suspension (Oct 2010)

Elog 3111 - June 2010, BS drag wiped.

  7915   Thu Jan 17 19:33:53 2013 ManasaUpdateAlignmentPRM oplev

 We had to work on redesigning the oplev layout in BSC when I found that the positions of the mirrors were clipping IPPOS and the green beam while updating the CAD layout.  

To avoid any clipping, the prm oplev beam is steered into the vacuum by an oplev mirror and out of vacuum through 3 steering mirrors. The table weights had to be moved to allow room for the oplev mirrors. Hence table had to be re-leveled. I will update the CAD drawing with the current position of the mirrors and will reconfirm that the new mirrors are not in the way of any of the beams. In-vac photos are updated in picasa.

  7914   Thu Jan 17 17:04:49 2013 SteveUpdateVACvac monitor screen changed

The COVAC_MONITOR.adl was changed. Ion pump labeled as disconnected means: ion pump controller power turned off and ion pump gate valve control connection disconnected.

RP2 roughing pump labeled disconnected means:  hardware  disconnected.

The actual operational, valve control screen has not been changed yet.

Attachment 1: screenmod.png
screenmod.png
  7913   Thu Jan 17 15:48:21 2013 JenneUpdateLockingPRM - Flat mirror cavity

 

 2" G&H mirror is installed on a DLC mount just in front of the BS.  I had to remove one of the 4 BS dog clamps, so we must put it back when we are finished with this test.

I aligned the G&H mirror such that the reflected beam is overlapped with the incident beam, and I aligned the PRM such that the regular REFL beam is retro-reflected.  This is the same as getting the beam bouncing off the PRM back to the G&H to be overlapped.

I then saw flashes of the cavity, when I held a card with a hole in the cavity, so the beam was going through a small aperture in the card, but I still saw flashes.  I was not able to see flashes on the IR card transmitted through the G&H mirror.

I also cannot see any flashes or scattered light on the face of PR2 camera.

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

Light is coming out of REFL on the AS table, but it's clipped somewhere....needs investigation/work before we can lock.

I also didn't see anything at the POP port with a card, but I'm hopeful that perhaps with a camera I'll see something.

  7912   Thu Jan 17 11:01:19 2013 JenneUpdateAlignmentYesterday's alignment work

[Jamie, Jenne, Manasa]

Yesterday's goal was to get the input beam centered on the PRM, the BS and ETMY simultaneously. 

Steve helped us remove the ETMY door first thing in the morning.  We then iterated with TT1, MMT1 and TT2 to try to get the beam centered on all the optics.  We were using MMT1 instead of TT1 for a while, so that we could keep TT1 in the center of its range, so that we had more range to use once we pump down.  Also, at one point, the beam was high on PRM, centered on BS, and high on ETMY, so Jamie poked PR3 a little bit.  This helped, although we closed up for lunch / group meeting soon after, so we didn't finalize any alignment stuff.

We decided to leave the rest of the full IFO alignment alone until after the PRM-flat test.

  7911   Thu Jan 17 01:27:54 2013 Tara(?)SummaryIOONoise budget for MC

Quote:

I missed the point.
Do you mean that we can measure the coating thermal noise of the ref cav at the 40m as the IMC is quiet enough?

 Yes, it should be. However, what I did was calculating thermal noise of MC. I'm not sure about the 40m IMC's actual noise level. The plot in the entry was taken from LLO's MC in 2003.

  7910   Thu Jan 17 00:17:31 2013 JenneUpdateLockingExpected PRC gains

Quote:

Why would we use such a bad optic in our recycling cavity? Is 1.5% the spec for these mirrors? Is this the requirement that Kiwamu calculated somehow? Did anyone confirm this measurement?

I can't believe that we'll have low noise performance in a RC where we dump so much power.

 Yeah, Koji mentioned in response to Raji's measurements several months ago that the LaserOptic mirros were pretty far out of spec. We should probably redo the measurement to confirm.

  7909   Wed Jan 16 20:27:16 2013 ranaUpdateLockingExpected PRC gains

Why would we use such a bad optic in our recycling cavity? Is 1.5% the spec for these mirrors? Is this the requirement that Kiwamu calculated somehow? Did anyone confirm this measurement?

I can't believe that we'll have low noise performance in a RC where we dump so much power.

  7908   Wed Jan 16 19:08:51 2013 KojiSummaryIOONoise budget for MC

I missed the point.
Do you mean that we can measure the coating thermal noise of the ref cav at the 40m as the IMC is quiet enough?

  7907   Wed Jan 16 18:58:08 2013 RijuUpdate Photodiode transimpedance

Today I have taken the reading for shot noise intercept current for the PDA255 - MC transmission RF PD. To do that I have put an incandescent bulb (JKL lamps, 222 bulbs, voltage and current rating 2.25V and 0.25A) in front of the PD and varied the current through it from 0A to 0.29A at 2.2V. I measured the corresponding DC voltage and took the noise data (4395A spectrum analyzer/ format noise, channel attenuation 0dB) through GPIB .

I will process the data and upload the result soon.

  7906   Wed Jan 16 18:52:49 2013 JenneUpdateLockingPRM - Flat mirror cavity plan

Game plan:

* Put 2" G&H mirror into BS chamber, in front of BS.

* Align it, lock cavity using an existing REFL PD.

* Align POP setup so I can use POP camera to take image of transmitted cavity mode, and actually take that image.

* Take image of face of PR2.

* Measure finesse of cavity using POP, or a Thorlabs PD at POP (looking at transmission through PR2) by scanning PRM, and infer cavity gain....compare with values in elog 7905.

* If time / inclination allow, take beam scan measurements of the REFL port.

I will not be able to do as was done in elog 6421 to look at the beam size at POP for non-resonating beams.  I expect ~0.1uW of light at POP in the non-resonant case:  100mW * 5.5% * 20ppm = 0.11microwatts.

 

  7905   Wed Jan 16 18:08:06 2013 JenneUpdateLockingExpected PRC gains

I was calculating the power recycling gains we expect for different versions of the PRC, and I am a little concerned that we aren't going to have much gain with the new LaserOptik mirrors.

I'm using

                       t_PRM^2

G =  -------------------------------------------

       (1 - r_PRM * r_PR2 * r_PR3 * r_end)^2

 

from eqn 11.20 in Siegman.

r_end is either the ITM (for a symmetric Michelson) or the flat mirror that we'll put in (for the PR-flat test case).

r = sqrt( R ) = sqrt( 1 - T ) for mirrors whose power transmission is the quoted value.

 

Some values: 

t_PRM^2 = T_PRM = 0.055   --------->   r_PRM = sqrt( 1 - 0.055 )

T_G&H = 20e-6   ---->   r_G&H = sqrt( 1 - 20e-6 )

T_LaserOptic = 0.015 (see elog 7624 where Raji measured this...1.5% was the best that she measured for P polarization.  Elog 7644 has more data, with 3.1% for 40deg AoI) -------> r_LasOpt = sqrt( 1 - 0.015 ) or sqrt( 1 - 0.031)

T_ITM = 0.014 -----------> r_ITM = sqrt( 1 - 0.014 )

 

Some calculations with 1.5% LaserOptik transmission:

G_PRC_2G&H = 45

G_PRC_G&H_LasOpt = 31

G_PRM_flatG&H = 51

With the 3% LaserOptik transmission:

G_PRC_G&H_LasOpt = 22

G_PRM_flatG&H = 30

More ideal case of just PRM, flat mirror (either ITM or G&H), ignoring the folding mirrors:

G_PRM_ITM = 45

G_PRM_flatG&H = 70

 

Punchline:

If the LaserOptik mirror has 1.5% transmission at ~45 degrees, the regular PRC expected gain goes down to 31, from 45 with both folding mirrors as G&Hs.

  7904   Wed Jan 16 10:57:37 2013 taraSummaryIOONoise budget for MC

I calculated thermal noise in mode cleaner (MC) mirrors and compared it with the measured MC noise. Thermal noise won't be a significant noise source for MC.

== Motivation==

 There is an idea of using MC and a refcav to measure coating thermal noise. One laser is frequency locked to MC, another laser is locked to an 8" refcav. Then the two transmitted beams are recombined so that we can readout the frequency noise. In this case, the transmitted beam from MC is a better reference (less frequency noise) than the beam from refcav. However, we need to make sure that we understand the noise sources, for example brownian noise, thermoelastic noise in both substrates and coatings, in MC more thoroughly.

==Calculation==

I used Rana's code for MC's technical noise sources from, svn. The same plot can be found in appendix C of his thesis. Then I added my calculation to the plot.  Jenne pointed me to 40m:2984 for the spot size and the cavity length. The spot radius on MC1 and MC3 is ~ 1.5mm, and ~3.4 mm@MC2, The round trip length is ~27m, thus the frequency fluctuation due to thermal noise is lower than that of refcav by 2-3 orders of magnitude. I calculated Brownian noise in coatings, Brownian noise in substrate, Thermoelastic noise in substrate. I assumed that the coatings are SiO2/Ta2O5, quarter stacks, coatings thickness for MC1/3 = 5um, for MC2 = 8um. The code can be found in the attachment.

mc_nb_TN.png

==result==

Total thermal noise on MC (Brownian + Thermoelastic on substrate and coatings of MC1-MC3) is plotted in dashed red. It is already below 10^-5 Hz/rtHz at ~20 Hz. This is sufficiently low compared to other noise sources. Beat signal from CTN measurement with 8" cavities is plotted in pink, the estimated coating brownian noise is plotted in a yellow strip. They are well above the measured MC noise between 100 Hz to a few kHz. Measuring coating thermal noise on 8" refcav seems plausible with this method. We can beat the two transmitted beams from IMC and refcav and readout the beat signal to extract the displacement noise of refcav. I'll discuss this with Koji if this is a good surf project.

 mc_nb_TN_2013_01_18.png

[the internal thermal noise in the original plotted is removed and replaced with the total thermal noise plot instead]

 note:I'm not sure about the current 40m MC configuration. The parameters used in this calculation are summarized in mcnoiseS2L1.m (in the svn page).

 

Attachment 2: mc_nb_TN.png
mc_nb_TN.png
Attachment 3: mc_nb_TN.fig
Attachment 4: MC_nb.m.zip
  7903   Wed Jan 16 08:29:45 2013 SteveUpdateSUSPRM damping restored

PRM oplev servo turned off.  OLPIT servo gain 0.15 and OLYAW  -0.3 set to ZERO.  PRM damping restored

  7902   Tue Jan 15 20:00:42 2013 ManasaUpdateAlignmentAdjustment of active TTs and input alignment

 

Just for reference! The changes made to the TT matrix in order to fix the polarity problem:

The old matrix values are mentioned in elog!

 

PIT    YAW                   New                   

Pit slider           |  -100   -100  |  UL  

     0               | -100    100   |  LL

Yaw slider           |  100   -100   |  UR

    0                |  100    100   |  LR

 

 

 

 

 

 

 

  7901   Tue Jan 15 19:26:35 2013 jamieUpdateAlignmentAdjustment of active TTs and input alignment

[Jamie, Manasa, Jenne]

We started by verifying that the tip-tilts were getting the correct signals at the correct coils, and were hanging properly without touching.

We started with TT2.  It was not hanging freely.  One of the coils was in much further than the others, and the mirror frame was basically sitting on the back side yaw dampers.  I backed out the coil to match the others, and backed off all of the dampers, both in back and the corner dampers on the front.

Once the mirror was freely suspended, we borrowed the BS oplev to verify that the mirror was hanging vertically.  I adjusted the adjustment screw on the bottom of the frame to make it level.  Once that was done, we verified our EPICS control.  We finally figured out that some of the coils have polarity flipped relative to the others, which is why we were seeing pitch as yaw and vice-versa.  At that point we were satisfied with how TT2 was hanging, and went back to TT1.

Given how hard it is to look at TT1, I just made sure all the dampers were backed out and touched the mirror frame to verify that it was freely swinging.  I leveled TT1 with the lower frame adjustment screw by looking at the spot position on MMT1.  Once it was level, we adjusted the EPICS biases in yaw to get it centered in yaw on MMT1.

I then adjusted the screws on MMT1 to get the beam centered at MMT2, and did the same at MMT2 to get the beam centered vertically at TT2.

I put the target at PRM and the double target at BS.  I loosened TT2 from it's base so that I could push it around a bit.  Once I had it in a reasonable position, with a beam coming out at PR3, I adjusted MMT1 to get the beam centered through the PRM target.  I went back and checked that we were still centered at MMT1.  We then adjusted the pitch and yaw of TT2 to get the transmitted beam through the BS targets as clear as possible.

At this point we stopped and closed up.  Tomorrow first thing AM we'll get our beams at the ETMs, try to finalize the input alignment, and see if we can do some in-air locking.

The plan is still to close up at the end of the week.

  7900   Tue Jan 15 01:41:40 2013 ranaUpdateSUSTT

 That seems like easily enough range; as long as we can put the TT into the middle of their range to start with we should be OK.

We should consider instrumenting the leakage transmission through all TT with a bare QPD on a stick. We can then use those sensors to monitor the spot positions within the input mirrors as well as the PRC / SRC.

  7899   Mon Jan 14 19:56:48 2013 ManasaUpdate TT

 

 [Manasa, Jenne]

The motion of the magnets (~1.5mm estimated by looking at the magnets moving) correspond to ~2deg. tilt of the mirror. This would mean almost 1.5m shift at the ETM end (~45m from the TT).

  7898   Mon Jan 14 15:15:04 2013 JenneUpdate TT

 

[Manasa, Jenne]

First plug in only one of the quadrupus cables, find out what coil it corresponds to according to screen, then plug in 2nd cable, don't test already-determined cable, but all other 3, find what cable it corresponds to according to the screen.  Repeat for other 2 cables.

TT2, confirmation:

C = LL, not UR, not UL, not LR

D = UL, not UR, not LR

A = LR, not UR

B = UR

 

---------------------------------------------------------------------------------------

After confirming that the correct quadrupus cables were plugged in to the correct coils, I suspected that our problems could be coming from a (or some) magnet(s) touching the inside of the OSEM.  We tested this a little bit, with the goal of finding the range of values where no magnets are touching.

All matrix values are either +1000 or -1000, so, with an example pitch slider value :

                       PIT    YAW

Pit slider           |  1000   1000  | --->  -22000 UL

     -22.2           | -1000   1000  | --->  +22000 LL

Yaw slider           |  1000  -1000  | --->  -22000 UR

    0                | -1000  -1000  | --->  +22000 LR

 

 

Trying some values for pitch, keeping yaw constant:

0 yaw, Pitch bias = 5 -> UR is touching on left side of its osem.

0 yaw, Pitch 0, UR is touching left side.

0 yaw, -1.2 pitch, UR just came off from touching left side.  More neg from here should be non-touching.  all others are fine.

0 yaw, -32.2 pitch, LR not quite touching right side of osem, but is close (much less than 1mm clearance).  UR fine. all others fine.

0 yaw, -22.2 pitch, all 4 are fine.

 

Trying some yaw values, keeping pitch constant:

1.  -22.2 pitch, -32 yaw, LR touching. UR touching.

2.  -22.2 pitch, -12 yaw, LR barely not touching, UR still touching.

3.  -22.2 pitch, 0 yaw, UR still touching.

4.  -22.2 pitch, 16 UR barely not touching.

5.  -22.2 pitch, 32, none touching.

6.  -22.2 pitch, 12, UR close, not touching.

7.  -22.2 pitch, 0, UR touching.

8.  -22.2 pitch, 32 (or 30?) UR came off.

9.  -22.2 pitch, -25, UR close

10.  -22.2 pitch, -32 UR touching.

11.  -22.2 pitch, -4 UR not touching.

12.  -22.2 pitch, 0 yaw, UR not touching.

 

Here is a graphical semi-representation for the yaw data:

 TT2_stuck_unstuck_14Jan2013.png

 

  7897   Mon Jan 14 12:08:39 2013 jamieUpdateAlignmentTT2 connections

Quote:

Quote:

Was the connection between the feedthrough (atmosphere side) and the connector on the optical table confirmed to be OK?

We had a similar situation for the TT1. We found that we were using the wrong feedthrough connector (see TT1 elog).

 The major problem that Manasa and I found was that we weren't getting voltage along the cable between the rack and the chamber (all out-of-vac stuff).  We used a function generator to put voltage across 2 pins, then a DMM to try to measure that voltage on the other end of the cable.  No go.  Jamie and I will look at it again today.

Everything was fine.  Apparently these guys just forgot that the cable from the rack to the chamber flips it's pins.  There was also a small problem with the patch cable from the coil driver that had flipped pins.  This was fixed.  The coil driver signals are now getting to the TTs.

Investigating why the pitch/yaw seems to be flipped...

  7896   Mon Jan 14 10:12:09 2013 JenneUpdateAlignmentTT2 connections

Quote:

Was the connection between the feedthrough (atmosphere side) and the connector on the optical table confirmed to be OK?

We had a similar situation for the TT1. We found that we were using the wrong feedthrough connector (see TT1 elog).

 The major problem that Manasa and I found was that we weren't getting voltage along the cable between the rack and the chamber (all out-of-vac stuff).  We used a function generator to put voltage across 2 pins, then a DMM to try to measure that voltage on the other end of the cable.  No go.  Jamie and I will look at it again today.

  7895   Mon Jan 14 09:08:37 2013 SteveUpdatePSLPMC is only ok short term

 

 The PMC PZT voltage slider seemed sticky.  First it would not do anything, than after moving slider back an forth a few times, it had a range of 60V and later it had full range and it locked

Attachment 1: pmc8d.png
pmc8d.png
  7894   Fri Jan 11 19:12:20 2013 KojiUpdateAlignmentTT2 connections

Was the connection between the feedthrough (atmosphere side) and the connector on the optical table confirmed to be OK?

We had a similar situation for the TT1. We found that we were using the wrong feedthrough connector (see TT1 elog).

  7893   Fri Jan 11 17:32:10 2013 ManasaUpdateAlignmentTT2 connections

 

 Manasa, Jenne

We started off to try and get TT2 working. We used the cables Jamie had already prepared while working on TT1 and used them to connect TT to the channels in 1Y3.

There were sma cable connectors already running between the channels 5-8 on the board to the UL,LL,UR and LR. Triggering the UL LL UR LR matrix on epics did not show any analog voltage at the output analog channels on the board. Talking to Jamie over phone, we inferred  that the  SMA cables that were already left connected corresponded to channels assigned for TT4 in epics.  So we set the connections right and could see analog voltage outputs corresponding to epics triggers.

We connected the ribbon cables running from the board to the TT. But changing pitch and yaw did not do anything to the TT2 mirror. We opened the BS door and checked if  the tt cables were connected to the post. We beeped the cable running from the board to TT (we also traced the cable's trail through the cable rack pile from 1Y3 to BSC). Using a function generator at the board end of the cable, we could not observe anything at the TT end of the cable.

We ran out of options on what can be done next and closed the doors. We hope Jamie can fix the problem once he returns next week.

  7892   Fri Jan 11 16:13:47 2013 JenneUpdateGeneralIFO status update - PMC fixed

Quote:

I came in this morning to see that the PMC was down. The PZT voltage had drifted to below 50V. I adjusted the FSS slow controls to 0V and PZT was back at 126V.

PMC and IMC could eventually be locked.

History of PZT voltage behaviour in dataviewer over the last 24 hours shows it has been drifting everytime after it has been fixed.

 FSS was saturating.  Fixed.

  7891   Fri Jan 11 11:07:04 2013 ManasaUpdateGeneralIFO status update - PMC problems

I came in this morning to see that the PMC was down. The PZT voltage had drifted to below 50V. I adjusted the FSS slow controls to 0V and PZT was back at 126V.

PMC and IMC could eventually be locked.

History of PZT voltage behaviour in dataviewer over the last 24 hours shows it has been drifting everytime after it has been fixed.

  7890   Thu Jan 10 15:30:33 2013 JenneUpdateAlignmentTT2 pins swapped, ITMX door open

 

[Bob, Manasa, Jenne]

We opened the ITMX heavy door.  Before getting too far, we realized that we had to do the fancy pin swapping before we can activate TT2.  So....

[Nic, Jenne]

We followed the instructions in elog 7869, and the associated Picasa album, and swapped the pins for the in-vac connector that will go to TT2.  Pretty easy, since the procedure was already well documented.

We then looked at the beam location on PR2, and the beam is ~2 inches up and to the left (as viewed from the front) from the center of the optic.  This is very easily correctable with the actuators, so we're leaving TT2 as it is.

  7889   Thu Jan 10 12:41:06 2013 ManasaConfiguration40m UpgradingEndtable upgrade for auxiliary green laser : wiki page

 

I have created a wiki page linked here with all details about the endtable upgrade.

The page has links to the new drawing for doubler post to hold the tilt-aligner that holds it. The Faradays will also be mounted similarly on tilt aligners placed on these posts. The bulk mounts will be made of aluminium similar to the colorful cylindrical mounts (images of which can be seen in the archived layouts on the wiki) that hold the He-Ne lasers and few faradays now.

  7888   Thu Jan 10 12:22:36 2013 JenneUpdateAlignmentTT matrix is funny

Quote:

* TT1 is in place, aligned so beam hits center of TT1, hits center of MMT1 (used pitch biases to finish pitch).

 I had asked Q to write this down on a piece of paper, but then I forgot to transcribe it into the elog....

The TT screen matrix, at least for TT1, is flipped or something.  When Eric moved the pit slider, the optic moved in yaw, and vice versa. 

We need to fix this, but for now, here's the situation when TT1 was pointed correctly at MMT1:

                       PIT    YAW

TT1 Pit slider     |  1000   1000  | --->   700 UL

     0             | -1000   1000  | --->   700 LL

TT1 Yaw slider     |  1000  -1000  | --->  -700 UR

    0.7            | -1000  -1000  | --->  -700 LR

 

The confusing thing is that Koji and I confirmed (by plugging in the correct cable to the correct sensor) that "UL" on the screen goes to the UL coil, and the same for the other 3 coils.  This needs investigation / fixing.

  7887   Wed Jan 9 19:32:24 2013 RijuUpdate Photodiode transimpedance

Summary:

Today I have tested the MC transmission-end RF photodiode PDA255 for transimpedance and dark noise using Jenne's Laser and AG4395A network/spectrum analyzer. The dark noise voltage distribution for the transmission and reflection PDs of MC and the analyzer has been compared.

Motivation:

I am to do the input mode cleaner cavity mode scan. The electronic and shot noise of the components used , particularly photodiode noise, will affect the peak position  of the modes, indicating the uncertainty in the measured frequencies of the modes. That will in turn give the uncertainty in the measured change of radius of curvature of the mirrors in presence of the laser beam, from which we will be able to calculate the uncertainty in the mirror-absorption  value.

Method:

For PD transimpedance measurement I used Jenne's laser along with AG4395 network analyzer. The RF out signal of AG4395A had been divided by splitter with one output of the splitter going to R channel of the network analyzer and the other to the laser. The splitted laser beams - splitted with beam splitter - fall on two photodiodes - one reference(Newfocus1617? PD, the DC and RF transimpedance values were taken from its datasheet ) and the other on PDA255. The outputs of these two photodiodes go to channel B and A respectively of the network analyzer. The measured transimpedance data had been collected using the GPIB connection. It had been ensured that the PD under test is not going to saturation, for that the source power level was kept to -40dBm. transimpedance measurements were compensated by the ratio of DC photocurrent.

For dark noise measurement the output of the PD was connected to the A channel of the AG4395A, when there was no light falling on it. The response is collected using GPIB. The attenuation of channel A was made 0dB. ( AG4395A was kept in Spectrum analyzer mode in Noise Format).

Results:

The plots corresponding to the measurements are attached.

Discussion:

The comparison for the dark noise voltage levels of the MC transmission PD (PDA255) with MC REFL PD has been made with analyzer dark noise voltage. It is shown in the attachment (I will upload the dark noise current comparison too....since the output darknoise depends on the gain of the circuit, it is important to divide this voltage spectra by transimpedances.)

Attachment 1: PDA255.pdf
PDA255.pdf
Attachment 2: PDA255_z.pdf
PDA255_z.pdf
Attachment 3: darknoiseVpda255.pdf
darknoiseVpda255.pdf
Attachment 4: darknoiseApda255.pdf
darknoiseApda255.pdf
Attachment 5: darknoise_comparison.pdf
darknoise_comparison.pdf
  7886   Wed Jan 9 18:59:01 2013 JenneUpdateAlignmentTT2 installed, PRM oplev layout changed

[Jenne, Manasa]

PZT2 was removed from the BS table, and packed away in a foil-lined plastic box.

PRM oplev path has been altered, including installation of a 3rd mirror, to accommodate TT2, which is larger than PZT2.

      * Unfortunately, PR3 is a few mm more north than is indicated in the CAD drawing, so I wasn't able to place the oplev mirrors exactly as Manasa indicated in elog 7815

      * We came up with a different layout. Photos were taken.  We will need to confirm that the IPPOS, AS, and GreenX beams all clear past the oplev mirrors, but by imagining straight lines between mirrors for those beams, I think we should be okay.  but we must confirm when we have real beams.

TT2 was installed, according to the placement in the diagram.   Dogged down just as TT1 - one dog for the riser, 3 dogs for the TT base which also squish the riser.  You should be able to see this in the photos. Without having installed the PRM target, it looks like the input beam is hitting pretty close to the PRM's center.  Tomorrow Jamie The Tall can install the PRM target for us so we can confirm.

 

Photos - I'm posting them on Picasa here.  The new camera, and the fact that you can rotate the viewfinder, is amazing for overhead in-chamber photos.  Seriously, it's much easier to take useful photos.  It's great.

 

Tomorrow:

We remove the ITMX door first thing.  If Steve isn't here, we'll ask Koji or Bob to help us with the crane. 

First thing on the alignment list is to finalize TT2's pointing.  Put a target in front of PR2, put on the PRM target, etc, etc.  We're basically back to the same alignment procedure as we've been doing the last few vents.

 

Item for meditation:

Do we trust ourselves, or do we want to think about installing a 'bathroom mirror' so we can see the face of PR3 while we are pumped down?

  7885   Wed Jan 9 13:34:34 2013 KojiSummaryGeneralThe projector lamp ended its life?

[Koji, Manasa]

- A new projector lamp installed.

- The old lamp lasted 8751 "equivalent lamp hours".

- The old lamp was found being shattered inside. It contains mercury.
So next time you hear the explosion sound of the lamp, establish the ventilation of the room and escape for an hour.

  7884   Tue Jan 8 18:10:41 2013 JenneUpdateSUSPRM, SRM, BS oplevs off

I don't know why (I'm just leaving the lab right now....) but BS, PRM, SRM all have no light on their oplev PDs. I have turned off the oplev servos for now, and will get back to them tomorrow, before redoing the BS table oplev layout.

  7883   Tue Jan 8 17:54:34 2013 JenneUpdateAlignmentRisers to bring TTs to correct beam height are in use

 

 [Jenne, EricQ, Nic, MattA]

* TT1 is in place, aligned so beam hits center of TT1, hits center of MMT1 (used pitch biases to finish pitch).

      * Riser installed, dogged down with 1 dog.

      * TT1 sitting on top of riser, 3 dogs holding TT to table, with riser squished in between.

* IOO table leveled.

      * Almost all of the weights on the IOO table were just sitting there, not screwed down!  One didn't even have a screw, 3 had screws, but they were totally loose.  2 of those screws were in as far as they could go, but they weren't holding the weight.  This means the screw was too long, and should have been replaced (which I did today).  Just because the existing screw was too long, doesn't mean it should be left as-is.  Everything in the chambers must be tightly clamped down, as soon as work on that item is complete!  Anyhow, after finalizing the leveling, I tightened down all of the weights on the IOO table.

* MMT1 tweaked so beam hits center of MMT2. 

* MMT2 tweaked so beam hits center of PZT2.

* Light access connector installed.

 

Sadface notes:

* I dropped a Class B golden-colored 3/16 allen key to the bottom of the IOO chamber.  I can't see it, but Nic thinks he might be able to see it with a mirror, from the BS chamber.  We should look for it when we look for the IR card that is still down there.

* We have an ant in the IOO chamber.  Unfortunately my hands were on the TT1 optic holder ring when I saw it, so I couldn't dash quickly enough to grab it.  I saw it run over the side of the table, and down, but looked under the table and couldn't find him.  Not so good, but I don't know what to do about it right now.  If anyone sees it, get it out please.

  7882   Tue Jan 8 15:28:41 2013 RijuUpdateElectronicsPhotodiode transimpedance

Quote:

Quote:

You have to correct this transimpedance ratio by correcting for the different levels of DC photocurrent in the two devices.

For the dark noise, you must always include a trace showing the noise of the measurements device (i.e. the analyzer noise must be less than the dark PD noise) with the same input attenuation setting.

 Hi,

The correction for different levels of DC photocurrent in the two devices had been taken care by one MATLAB code, the code that originally was made by Koji.

The analyzer noise I had not recorded; today I am going to record it.

Riju

 Here is the data for AG4395A network/spectrum analyzer noise data. I collected the data by putting 50ohm terminator on channel A with same input attenuation setting (0dB attenuation).

Attachment 1: analyzernoiseV.pdf
analyzernoiseV.pdf
  7881   Tue Jan 8 14:07:04 2013 RijuUpdateElectronicsPhotodiode transimpedance

Quote:

You have to correct this transimpedance ratio by correcting for the different levels of DC photocurrent in the two devices.

For the dark noise, you must always include a trace showing the noise of the measurements device (i.e. the analyzer noise must be less than the dark PD noise) with the same input attenuation setting.

 Hi,

The correction for different levels of DC photocurrent in the two devices had been taken care by one MATLAB code, the code that originally was made by Koji.

The analyzer noise I had not recorded; today I am going to record it.

Riju

  7880   Tue Jan 8 14:01:21 2013 RijuUpdate Photodiode transimpedance

 Here I upload the plots corresponding to my last day's measurements.

 

Attachment 1: TFreflpd.pdf
TFreflpd.pdf
Attachment 2: REFL_z.pdf
REFL_z.pdf
Attachment 3: darknoiseVreflpd.pdf
darknoiseVreflpd.pdf
Attachment 4: darknoiseAreflpd.pdf
darknoiseAreflpd.pdf
  7879   Mon Jan 7 19:23:19 2013 ranaUpdateElectronicsJamie's 1811 PS from 1998

  1. Front Panel switch supplies power, but does not light up - its unsafe as is. Needs new switch.
  2. Output has current limiting (which is nice) and schematic inside the box (which is very nice).
  3. Output voltage is not filtered or regulated ? LM7812 / 7912 would do the trick - or pick a PS with 18V outputs to reg down to 15 V.
  4. Box needs rubber feet.
  5. Overall B-

Also, we still need to get a 32GB SD card for the new camera. It only has an 8GB one.

Attachment 1: SOLA-1021-1026.pdf
SOLA-1021-1026.pdf SOLA-1021-1026.pdf SOLA-1021-1026.pdf SOLA-1021-1026.pdf SOLA-1021-1026.pdf SOLA-1021-1026.pdf
Attachment 2: jamie98.jpg
jamie98.jpg
  7878   Mon Jan 7 16:45:30 2013 JenneUpdateAlignmentRisers to bring TTs to correct beam height are wrong

[Jenne, with backup from Koji and Steve]

Short version:

TT1 was installed without a riser, optic is too low, riser we have doesn't fit, cannot proceed with alignment.  Sadface.

Long version:

I had gotten to the point of checking that the beam coming out of the Faraday was hitting the center of TT1, when I realized that we had forgotten to install the risers.  The TTs are designed for 4" beam height, but we have a beam height of 5.5" in-vac.  This means that the beam out of the Faraday was hitting the top of the optic / the optic holder.

Steve showed me where all of the active TT equipment is stored (down the X arm, almost all the way to the flow bench...there is a plastic tub full of baked items (individually wrapped and bagged)), and I got one of the 1.5" risers.

Upon opening the riser package, and comparing it with the base plate of the active tip tilt, the screw holes don't match!

TT1_7Jan2013_BasePlateAndRiserHoleMismatch.JPG

It looks like for the passive tip tilts, we had holes machined at the far corners of the base plate, then had these risers made.  You can see in the photo of SR3 below that the original holes are there, but we are using 1/4-20 holes at the far corners of the base plate.

SR3_7Jan2013_ExtraHoleMachinedAtCorner.JPG

Unfortunately, without checking the base plate, I had asked Steve to get 4 more of the same risers we used for the passive tip tilts.  So, now the base plate holes and the riser holes don't match up.  In a perfect world, we would have installed the risers on the TTs as soon as they were baked and ready, and would have discovered this a while ago....but we don't live in that world.

 The reason we had originally chosen to put the new 1/4-20 holes on the corners of the passive tip tilts was so that when we tightened the screws, we wouldn't bend the base plate, due to the groove at the bottom of the base plate being directly under the screws.  Since the new aLIGO TT base plates have the groove underneath going the opposite direction, we didn't need to move the holes to the corners.

Also, you can't really see this from the photos, but the active TT base plate is slightly longer (in the beamline direction) than the riser, but only by a little bit.  Koji is currently trying to measure by how much from the CAD drawings.

Also, also, because of the way TT1 will hang off the table, I'm concerned about the underneath groove on the riser being the direction it is.  I'm concerned that the grooved part will be what wants to touch down on the back corner of the table, such that either the TT is insufficiently supported, or it is tilting backwards.  Neither of these will be acceptable.

I propose that we re-make the risers quickly.  We will have the holes match the active TT base plate, the size of the riser should match the size of the active TT base plate, and the underneath groove should be perpendicular to the way it is in the current version.

  7877   Mon Jan 7 00:08:16 2013 JenneHowToAlignmentIn-vac plan

List of things to do, in order:

* Remove BS heavy door.  Steve, please remove the BS door as soon as you have enough people to do so.  I will be a little late, since I have a dentist appointment, but please don't wait for me.  Jamie and Manasa can help you.  Put on a light door.

* Remove MC light doors, make aluminum foil tube (not light access connector, yet).

* Open laser shutter, lock PMC. (Required slight tweaking of input steering.) Confirm power level into vacuum <100mW.

* Lock MC and check spot positions of MC (quickly.  this shouldn't take all day, hopefully).

-------------------------------  End of work for Monday.  See following elog ------------------------------------------------

 

* Move TT1 to be as close as possible to the location indicated on the diagram, then align it. 

         * Make sure beam out of Faraday is hitting the center of the optic.

         * Make sure beam reflected off of TT1 hits center of PZT2.  Only use actuators for the final alignment, then confirm that they aren't close to the edge of their ranges.

         * Lock down TT1 with dog clamps.

* Put light access connector on MC.

* Swap PZT2 out with TT2.  Should be at correct spot, according to diagram, and beam should be hitting center of optic.  Alignment only to the ~few degree point here.

* Re-level BS table.

* Fix oplevs that need fixing.  (Manasa should have the plan on one of the diagrams).

* Put target on PRM cage.

* Align TT2 so that beam goes through PRM target.

* Open ITMX heavy door. (Probably Tuesday morning).

* Place freestanding target in front of PR2.  Ensure TT2 is aligned to go through PRM target, and hit center of PR2.  Again, save actuators for fine-tuning.

At this point, I think we should (temporarily) install one of the G&H mirrors as a flat mirror facing the PRM, and see if we can lock that cavity using REFL.  We will want to have already created a model for this case, to compare our observations to.  Or we could align the full PRMI, and try to lock that in air.

 

 

  7876   Fri Jan 4 15:11:28 2013 JenneUpdate TT

[Jenne, Koji]

D - UL

B - UR

A - LR

C - LL

The sensor card on the bottom of the chamber was not salvaged yet.

  7875   Fri Jan 4 13:23:10 2013 ranaUpdateElectronicsPhotodiode transimpedance

You have to correct this transimpedance ratio by correcting for the different levels of DC photocurrent in the two devices.

For the dark noise, you must always include a trace showing the noise of the measurements device (i.e. the analyzer noise must be less than the dark PD noise) with the same input attenuation setting.

  7874   Thu Jan 3 20:34:43 2013 RijuUpdate Photodiode transimpedance

Today I have measured the transimpedance and dark-noise of the MC-REFL PD.

For transimpedance measurement I first collected the data of the reference Newfocus PD connecting it at channel B of Network-analyzer using the set-up of Jenne's laser. The data for the MC-REFL PD had been collected by connecting it to the A channel of Network Analyzer. To do that I shifted the Jenne's Laser to the table of MC-REFL PD, I moved the laser output on the table and fixed a lens and a mirror on the table. Taking the ratio of the two sets of datas I got the required trans-impedance.

Dark-noise readings were taken keeping the laser off.

I will upload the corresponding plots tomorrow.

  7873   Thu Jan 3 19:19:59 2013 ranaHowToElectronicscable racks

Today I found 3 power cables in the orange Pomona cable tray, put in so that the cables were damaged and therefore dangerous.

Please think about what you are doing before doing it. Damaging these things because your are in a hurry or frustrated will just waste our time and damage our interferometer.

For reference, we only use the thick blue Pomona racks for power cables. We use the orange and black ones for thinner cables. Pay attention and keep the cables organized.

Cable Rack Selection

 

  7872   Wed Jan 2 15:33:23 2013 JenneHowToLockingWe should retry in-air locking

Immediate things to do include finishing installation of new TTs and re-routing of oplev paths in the BS chamber, but after all that, we should retry in-air locking.

The last time we (I) tried in-air locking, MICH wouldn't lock since there was only ~ 6uW of light on AS55 (see elog 7355).  That was before we increased the power into the MC by a factor of 10 (see elog 7410), so we should have tens of microwatts on the PD now.  At that time, we could barely see some PDH signal hidden in the noise of the PD, so with a factor of 10 optical gain, we should be able to lock MICH.

REFL should also have plenty of power - about 1.5 times the power incident on the PRM, so we should be able to lock PRCL. 

Even if we put a flat G&H mirror after the PRM to make a mini-cavity, and we lose power due to poor mode matching, we'll still have plenty of power at the REFL port to lock the mini-cavity.

For reference, I calculate that at full power, POX and POY see ~13uW when the arms are locked.

 

POX/POY power =  [  (P_inc on ITM) + (P_circ in arm)*(T_itm)  ] * (pickoff fraction of ITM ~ 100ppm)

REFL power = (P_inc on PRM) + (P_circ in PRCL)*(T_prm)     =~ 1.5*(P_inc on PRM)

  7871   Wed Jan 2 06:52:50 2013 KojiUpdateSUSTT in vac DB25 pin swapping

[Koji, Rana, Nic, Steve]

I recalled that we used an optical lever to check if the TT is moving or not.
We used a laser pointer on a tripod, which was prepared by Steve.

I should also note that we stepped back the eddy current dampers from the magnets
in order to enhance the motion of the suspension. They should be restored in the end.

The mini-D connectors on the OSEMs are loosened so that we can plug the cables in.
This requires a specific metric allen key
that is stored in a clean tool box with an aluminum foil.

  7870   Fri Dec 21 19:49:39 2012 RijuUpdate Photodiode transimpedance

I have repeated the transimpedance measurement of PDA10CF. Also made the dark current noise measurement by connecting the PDA10CF output to the A channel of network analyzer.  The results are as follows. I I started to take the reading for shot noise intercept current using a light bulb in front of the PD, changing the current through the bulb, but at higher current the bulb filament got broken, so the experiment is incomplete.

Attachment 1: PDA10CFrepeat.pdf
PDA10CFrepeat.pdf
Attachment 2: darknoiseVpda10cf.pdf
darknoiseVpda10cf.pdf
Attachment 3: darknoiseApda10cf.pdf
darknoiseApda10cf.pdf
Attachment 4: PDA10CF_z.pdf
PDA10CF_z.pdf
  7869   Fri Dec 21 16:50:30 2012 RanaUpdateSUSTT in vac DB25 pin swapping

[Koji, Rana, Nic, Steve]

We went to the 25-pin D cable which connects to the TT1 quadropus and succeeded eventually in swapping pins 12/24 into the 13/25 positions.

  1. The D-sub connector is a custom made LIGO part and so it doesn't at all work to use the standard pin extractor tools to move the pins out; we should have investigated this before spending all this time poking at and possibly damaging the existing connector.
  2. To move the pins, we have to partially dis-assemble the connector and fish the pins/wires through the appropriate holes. Unfortunately, the design is such that we nearly lose all of the pins when trying to do this. Pictures describe the story better than words.
  3. After the swap we tried to test the TT, but again wasted some time because the vac feedthrough was incorrectly labeled. The 25-pin feedthrough labeled as "PZT1" does not, in fact, connect to the TT. Instead, its the one slightly above it that is labeled "Pico". I have moved the PZT1 sticker up to match the actual connector. In order to discover this, we beeped through several stages of the coil driver, cable system. WE need to order some in-line D-sub breakouts for 25pin, 37pin, and 9pin which are similar to the ones we have now for 15pin. These are better than the green terminal block breakouts.
  4. After this, we were able to see the TT move, but elected to leave the final piece of the work (determining which microD goes with which coil) to when Jamie gets back.
  5. The TT screen is not good: it needs to be just like the usual sus screen so that we can put in offsets, excitations, etc. Perhaps also the ASC-TT screen can link to the TT:SUS screens. We can just copy the eLIGO TT screens to get going.
ELOG V3.1.3-